OpenStudy (anonymous):

Find an equation of the line containing the given pair of points. Write your final answer in slope-intercept form. (5, 0) and (1, -3) Is it y=-3/4 x - 15/4 or y= 3/4 x - 15/4? > _ >''

OpenStudy (anonymous):

\[m=\frac{y2-y1}{x2-x1}\]

OpenStudy (anonymous):

m = -3-0 / 1-5 = -3/-4 = 3/4 now using find an equation of a line formula and one point (5,0) y2-y1=m(x2-x1) y-0 = 3/4(x-5) y = 3/4x - 15/4

OpenStudy (anonymous):

so ur second one.

OpenStudy (anonymous):

I get really confused whether it's negative or not > __ >'' Thank you <3 c:

OpenStudy (anonymous):

no problem just remember wen you are writing it out to put -- or -(-4) or some sort of indication to remind you that -- equals a positive. :)